OpenStudy (anonymous):

Does anyone know how to solve backwards? Write a trigonometric equation with solutions 5pi/6 and 7pi/6. Check by solving your equation.

OpenStudy (anonymous):

you have to look at your unit circle for that. I suggest that for future studies you memorize the thing since precal is pretty much all about that. i'll explain.

OpenStudy (rob1525):

use the unit circle and the law of sin and cosine. Yes, and like Aero_Exis writes MEMORIZE IT!

OpenStudy (anonymous):

on your unit circle you can see where 5pi/6 and 7pi/6 are in quadrant 2 and 3

OpenStudy (anonymous):

so basically you have the cos or sin that can equal those two values in those two quadrants.

OpenStudy (anonymous):

you may use either one. the cos(-sqrt3/2) can get you to 5pi/6

OpenStudy (anonymous):

as shown on the unit circle. and in effect that equation also gets you 7pi/6 as well because sqrt3/2 is negative in two areas on the unit circle

OpenStudy (anonymous):

at 150 degrees and at 210

OpenStudy (anonymous):

I did 3tan(x)^2 - 1=0

OpenStudy (anonymous):

i can say that you won't see anything to the power of 2 in the unit circle

OpenStudy (amistre64):

trig((5pi/6-x)(7pi/6-x)) = trig(0)